# Diffusion Variational Inference: Diffusion Models as Expressive Variational Posteriors

### v1

@@ 3.1.1 Denoising Diffusion Variational Inference (DDVI)
$\mathcal{L}$ $=\underbrace{\mathbb{E}_{q_{\bm{\phi}}(\mathbf{y},\mathbf{z}|\mathbf{x})}[\log p_{\bm{\theta}}(\mathbf{x}|\mathbf{z})]}_{\text{wake / recons. term $\mathcal{L}_{\text{rec}}(\mathbf{x},{\bm{\theta}},{\bm{\phi}})$}}-\underbrace{D_{\mathrm{KL}}(q_{\bm{\phi}}(\mathbf{y},\mathbf{z}|\mathbf{x})||p_{\bm{\theta}}(\mathbf{y},\mathbf{z}))}_{\text{prior regularization term $\mathcal{L}_{\text{reg}}(\mathbf{x},{\bm{\theta}},{\bm{\phi}})$}}-\underbrace{\mathbb{E}_{p_{\bm{\theta}}(\mathbf{x})}[D_{\mathrm{KL}}(p_{\bm{\theta}}(\mathbf{z}|\mathbf{x})||q_{\bm{\phi}}(\mathbf{z}|\mathbf{x}))]}_{\text{sleep term $\mathcal{L}_{\text{sleep}}({\bm{\phi}})$}}$ (3)
Observe that this objective is the ELBO in Equation 2 augmented with an additional regularizer $\mathcal{L}_{\text{sleep}}({\bm{\phi}})$ that is inspired by the sleep phase of the wake-sleep algorithm. This term consists of the reverse KL divergence $D_{\mathrm{KL}}(p_{\bm{\theta}}(\mathbf{z}|\mathbf{x})||q_{\bm{\phi}}(\mathbf{z}|\mathbf{x}))$. Section 3.1.2 shows that this divergence can be optimized via a diffusion-like training procedure. As in wake-sleep, we optimize $\mathcal{L}_{\text{sleep}}$ over ${\bm{\phi}}$ only.
Objective (3) poses restrictions on $p_{\bm{\theta}}$ and $r$. The prior $p_{\bm{\theta}}(\mathbf{y},\mathbf{z})$ needs to have a tractable density, although we will also present approximations and experimental results with implicit sample-based priors. Additionally, $q_{\bm{\phi}}(\mathbf{y},\mathbf{z}|\mathbf{x})$ must feature tractable entropy and sampling.
Lastly, note that $\mathcal{L}$ lower bounds the marginal log-likelihood $\log p_{\bm{\theta}}(\mathbf{x})$. When $q_{\bm{\phi}}(\mathbf{y},\mathbf{z}|\mathbf{x})$ equals the true posterior, this bound is tight, since the ELBO is tight, and the sleep term also equals zero.
3.1.2 Optimization Using Wake-Sleep in Latent Space
Next, we introduce optimization algorithms for diffusion variational inference. Maximizing $\mathcal{L}(x,{\bm{\theta}},{\bm{\phi}})$ involves optimizing the sleep term $\mathcal{L}_{\text{sleep}}({\bm{\phi}})$. This optimization is tractable since:

### v3

@@ 3.2 The Wake-Sleep Regularized ELBO
$\log p_{\bm{\theta}}(\mathbf{x})\geq\underbrace{\mathbb{E}_{q_{\bm{\phi}}(\mathbf{y},\mathbf{z}|\mathbf{x})}[\log p_{\bm{\theta}}(\mathbf{x}|\mathbf{z})]}_{\text{wake / recons. term $\mathcal{L}_{\text{rec}}(\mathbf{x},{\bm{\theta}},{\bm{\phi}})$}}\underbrace{-D_{\mathrm{KL}}(q_{\bm{\phi}}(\mathbf{y},\mathbf{z}|\mathbf{x})||r(\mathbf{y}|\mathbf{x},\mathbf{z})p(\mathbf{z}))}_{\text{prior regularization term $\mathcal{L}_{\text{reg}}(\mathbf{x},{\bm{\theta}},{\bm{\phi}})$}}\underbrace{-\mathbb{E}_{p_{\bm{\theta}}(\mathbf{x})}[D_{\mathrm{KL}}(p_{\bm{\theta}}(\mathbf{z}|\mathbf{x})||q_{\bm{\phi}}(\mathbf{z}|\mathbf{x}))]}_{\text{sleep term $\mathcal{L}_{\text{sleep}}({\bm{\phi}})$}}$ (8)
The optimization of the regularizer $\mathcal{L}_{\text{sleep}}({\bm{\phi}})$ is similar to the sleep phase of the wake-sleep algorithm, and closely resembles diffusion model training (see below). As in wake-sleep, $\mathcal{L}_{\text{sleep}}$ is optimized over ${\bm{\phi}}$ only, and the $\mathbf{x}$ are sampled from the model.
3.3 Optimizing the Regularized ELBO
Computing $\mathcal{L}_{\text{sleep}}({\bm{\phi}})$ still involves intractable distributions $p_{\bm{\theta}}(\mathbf{z}|\mathbf{x}),q_{\bm{\phi}}(\mathbf{z}|\mathbf{x})$. To optimize $\mathcal{L}_{\text{sleep}}({\bm{\phi}})$, we introduce another lower bound $\mathcal{L}_{\text{diff}}({\bm{\phi}})$, which we call the denoising diffusion loss (for reasons that will become apparent shortly):
$\mathcal{L}_{\text{sleep}}({\bm{\phi}})$ $=-\mathbb{E}_{p_{\bm{\theta}}(\mathbf{x})}[D_{\mathrm{KL}}(p_{\bm{\theta}}(\mathbf{z}|\mathbf{x})||q_{\bm{\phi}}(\mathbf{z}|\mathbf{x}))]=\mathbb{E}_{p_{\bm{\theta}}(\mathbf{x},\mathbf{z})}[\log{q_{\bm{\phi}}(\mathbf{z}|\mathbf{x})}]+\bar{H}(p_{\bm{\theta}})$ (9)
$\geq\mathbb{E}_{p_{\bm{\theta}}(\mathbf{x},\mathbf{z})}[\mathbb{E}_{r}[\log\frac{q_{\bm{\phi}}(\mathbf{y},\mathbf{z}|\mathbf{x})}{r(\mathbf{y}|\mathbf{z},\mathbf{x})}]+\bar{H}(p_{\bm{\theta}})=\mathcal{L}_{\text{diff}}({\bm{\phi}})$ (10)
In Equation 10, we applied the ELBO with $r(\mathbf{y}|\mathbf{z},\mathbf{x})$ playing the role of the variational posterior over the latent $\mathbf{y}$ in $q_{\bm{\phi}}$; $\bar{H}(p_{\bm{\theta}})$ is the expected conditional entropy of $p_{\bm{\theta}}(\mathbf{z}|\mathbf{x})$, a constant that does not depend on ${\bm{\phi}}$.
We define the DDVI learning objective $\mathcal{L}(\mathbf{x},{\bm{\theta}},{\bm{\phi}})$ as the sum of the aforementioned terms:
$\mathcal{L}(\mathbf{x},{\bm{\theta}},{\bm{\phi}})=\mathcal{L}_{\text{rec}}(\mathbf{x},{\bm{\theta}},{\bm{\phi}})+\mathcal{L}_{\text{reg}}(\mathbf{x},{\bm{\theta}},{\bm{\phi}})+\mathcal{L}_{\text{diff}}({\bm{\phi}})$ (11)
Terms $\mathcal{L}_{\text{reg}}$ and $\mathcal{L}_{\text{diff}}$ may be weighted by hyper-parameters $\beta_{\text{reg}},\beta_{\text{diff}}>0$, as in the $\beta$-VAE framework. In our experiments, assume $\beta_{\text{reg}}=\beta_{\text{diff}}=1$ unless otherwise specified. Note that since $\mathcal{L}_{\text{diff}}\leq\mathcal{L}_{\text{sleep}}\leq 0$, $\mathcal{L}(\mathbf{x},{\bm{\theta}},{\bm{\phi}})$ is a valid lower bound on $\log p_{\bm{\theta}}(\mathbf{x})$ that is tight when $q_{\bm{\phi}}(\mathbf{z}|\mathbf{x})=p_{\bm{\theta}}(\mathbf{z}|\mathbf{x})$.
Wake-Sleep Optimization
We optimize our bound on $\mathcal{L}(\mathbf{x},{\bm{\theta}},{\bm{\phi}})$ using gradient descent by alternating between ELBO optimization and taking sleep steps (see Section 5). Note that by maximizing $\mathcal{L}_{\text{diff}}$, we fit $q_{\bm{\phi}}(\mathbf{z}|\mathbf{x})$ to $p_{\bm{\theta}}(\mathbf{z}|\mathbf{x})$ via the forward KL divergence; similarly, by optimizing $\mathcal{L}_{\text{rec}}+\mathcal{L}_{\text{reg}}$ (the ELBO), we fit $q_{\bm{\phi}}(\mathbf{z}|\mathbf{x})$ to $p_{\bm{\theta}}(\mathbf{z}|\mathbf{x})$ via the reverse KL divergence. Thus, optimizing $\mathcal{L}(\mathbf{x},{\bm{\theta}},{\bm{\phi}})$ encourages $q_{\bm{\phi}}(\mathbf{z}|\mathbf{x})$ to approximate $p_{\bm{\theta}}(\mathbf{z}|\mathbf{x})$, and when the two are equal, the bound on $\log p_{\bm{\theta}}(\mathbf{x})$ is tight.
Simplifying Wake-Sleep
We also consider a light-weight algorithm, in which $r(\mathbf{y}|\mathbf{z})$ and $q_{\bm{\phi}}(\mathbf{z}|\mathbf{y})$ do not depend on $\mathbf{x}$. In this case, $\mathcal{L}_{\text{diff}}$ requires only sampling from $p(\mathbf{z})$, and the entire loss $\mathcal{L}$ can be optimized end-to-end using gradient descent. This algorithm is a simpler (there is no separate sleep phase); however, $q_{\bm{\phi}}(\mathbf{z}|\mathbf{x})$ may not perfectly approximate $p_{\bm{\theta}}(\mathbf{z}|\mathbf{x})$, hence $\mathcal{L}$ may no longer be a tight bound. We report results in Appendix B.
3.4 Combining The Wake-Sleep Regularized ELBO With Diffusion Models
Lastly, we can further simplify $\mathcal{L}_{\text{diff}}$ by leveraging the Markov structure of the forward and reverse processes $r,q$. Recall that each $\mathbf{y}=(\mathbf{y}_{1},\mathbf{y}_{2},...,\mathbf{y}_{T})$ can be a vector of $T$ latents, which we also denote as $\mathbf{y}_{1:T}$, and that $r(\mathbf{y}_{1:T}|\mathbf{z})=\prod_{t=1}^{T}r(\mathbf{y}_{t}|\mathbf{y}_{t-1})$, where $\mathbf{y}_{0}=\mathbf{z}$ and the $\mathbf{y}_{1:T}$ are increasingly noised versions of $\mathbf{y}_{0}$. Similarly, $q_{\bm{\phi}}(\mathbf{y},\mathbf{z}|\mathbf{x})$ is an approximate reverse diffusion process $q_{\bm{\phi}}(\mathbf{y},\mathbf{z}|\mathbf{x})=q_{\bm{\phi}}(\mathbf{y}_{0:T}|\mathbf{x})=q_{\bm{\phi}}(\mathbf{y}_{T}|\mathbf{x})\prod_{t=1}^{T}q_{\bm{\phi}}(\mathbf{y}_{t-1}|\mathbf{y}_{t},\mathbf{x})$ that is trained to reverse $r$.
In order to fit $q_{\bm{\phi}}$, we form a lower bound $\mathcal{L}_{\text{diff}}(\mathbf{x},{\bm{\phi}})$ on the sleep term $\mathbb{E}_{p(\mathbf{x},\mathbf{z})}\log{q_{\bm{\phi}}(\mathbf{z}|\mathbf{x})}$ in Equation 9. This bound is identical to the ELBO of a diffusion model, and has the same derivation.